OpenStudy (anonymous):

How do I find the values of theta if cot theta= √3/3?

OpenStudy (anonymous):

\[\cot \theta =\frac{ \sqrt{3} }{ 3 }\]

OpenStudy (anonymous):

By definition, cotangent is 1 over tangent. So, \[\tan x = 1/\cot x = \frac{ 3 }{ \sqrt{3} }\] From here, you could plug in this value to your calculator and find the arctangent.

OpenStudy (anonymous):

I got 1.732 but how do I get the values in radians or degrees?

OpenStudy (anonymous):

1.732 is the value of our tangent. So, we need to find the arctan of this value. You can do this on most scientific calculators. You could also use google's calculator if you'd like: http://goo.gl/QJdFx

OpenStudy (anonymous):

I tried the second question \[\cos \theta =\ \frac{ 1 }{ 2 }\] I got pi/6 is that correct?

OpenStudy (anonymous):

You're close, thats actually the arcsin of 1/2 and not the arccos. If it said \[\sin \Theta = \frac{ 1 }{ 2 }\] then you'd be correct. But you have to evaluate arccosine(1/2) which you can plug into google if you'd like. The answer you should get is pi/3 or 60 degrees.

OpenStudy (anonymous):

Okay I see, thank you

OpenStudy (anonymous):

 I'm not exactly sure how to solve this